Father age is 4 times as much as the sum of the ages of his three children but 6 years after his age will be only double the sum of his son age . Find his age
Answer by mananth(16946)

Let sum of ages of children be x
Father's age = 4x
After 6 years father's age = 4x+6
sum of children's age = x+3*6=x+18 ( each child's age increases by 6 so three children's age will increase by 18)
4x+6=x+18
3x=12
x=12
Father's age = 12*4=48 years
m.ananth@hotmail.ca